[Brahms-l] Shift report 20040326 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Fri Mar 26 2004 - 07:37:17 EST
------------------------------------------------------------
Date: Fri Mar 26 07:37:17 2004, Shift: 00:00-08:00
Supervisor: EBJ, with: 
Log book#15 pages 99 - 99

Shift Summary:
------------------------------------------------------------
1.  WE received 4 hours of beam from midnight to 8am.  

2.  I took data at FS 3 deg. 1/2 field B Pol (~500k Trig 2) and MRS 40 deg @ 1050 A Pol ( ~150k Trig3)

3.  I noticed that the DVM distributions in supermon look wierd.  See the Elog.
  The BERT page doesn't seem to be working.  I called MCR twice about it, and they said once an expert arrives for work, he will take a look at our problem.  

------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Mar 26 07:37:27 2004

This archive was generated by hypermail 2.1.8 : Fri Mar 26 2004 - 07:37:47 EST